OpenStudy (anonymous):

x-14/x-1=4/y can anyone help me solve please

OpenStudy (zehanz):

Is this what you mean:\[\frac{ x-14 }{x-1 }=\frac{ 4 }{ y }\]Solving it for y would go like this: \[y(x-14)=4(x-1)\]so\[y=\frac{ 4(x-1) }{ x-14 }\]If the equation is different, please try to use the equation editor or use brackets to avoid ambiguity. In the calculation above I used this rule: \[\frac{ a }{ b }=\frac{ c }{ d }\rightarrow ad=bc\]

OpenStudy (anonymous):



OpenStudy (zehanz):

OK, that is an equation in x (y is gone?). You can multiply everything with x-1 to get rid of the fractions:\[x(x-1)-14=4(x-1)-2x\] If you multiply out all the brackets and put the terms in the right order (first x², then the term with x, then a constant), you'll see that it is a 2nd degree equation. Do you know how to solve that?

OpenStudy (anonymous):

do I need to get x by itself

OpenStudy (anonymous):

or no figure outhe two x values

OpenStudy (zehanz):

If you solve the 2nd deg eq. you will get 2 solutions for x.

OpenStudy (zehanz):

\[x^2-x-14=4x-4-2x\]\[x^2-3x-10=0\]You can solve this by factoring.

OpenStudy (anonymous):

ok x=-2 or x=5

OpenStudy (zehanz):

You solved it!

OpenStudy (anonymous):

THANK YOU